# On two maximally entangled couples

### v1

@@ Discussion
Finally, we should mention a proof that we skipped: The proof by Grassl, Beth, and Pellizzari shows that no $(\!(3,2,3)\!)$ exists. As a consequence of, also no $(\!(4,1,3)\!)$. This proof is rather intricate [GBP97, Theorem 5], relying on a sequence of arguments involving the Knill-Laflamme conditions, and the fact [GBP97, Lemma 3] that for every two–dimensional subspace of $\mathds{C}^{2}\otimes\mathds{C}^{2}$ there exists a basis that contains at least one product state. This leads to the conclusion that a $(\!(3,2,3)\!)$ code can be factored into a tensor product, which contradicts that there does not exist a non-trivial code of length two.
@@ Appendix A Three-qubit GHZ states
The proof is carried out by explicit construction of the local unitaries required to transform a given state to the standard GHZ form. Consider a three-qubit state $\mathinner{|\psi_{ABC}\rangle}$ with three maximally mixed 1-RDM. We start with a Schmidt decomposition with respect to the bipartite cut that contains qubit $A$,
@@
where $a$ and $b$ are complex numbers. This state can be substituted in $\mathinner{|\psi_{ABC}\rangle}$. If we now compute the reduced states $\rho_{B}$ and $\rho_{C}$ we find that the off-diagonal elements of $\rho_{B}$ up to complex conjugation - are proportional to
(47) $-\sqrt{\lambda}\ a^{*}\ +\ b\ \sqrt{1-\lambda}\ ,$
@@
For $\lambda\neq\frac{1}{2}$ and $a,b\neq 0$ vanishing off-diagonal elements of the 1-RDMs in Eq. (47) and (48) is not possible. Therefore it must hold that $a=b=0$. Hence for $\lambda\neq\frac{1}{2}$ we obtain for $\psi_{ABC}$,
@@
[…]
@@
If instead $\lambda=\frac{1}{2}$ we have $b=a^{*}$ then $\mathinner{|u_{1}\rangle}$ becomes
@@
$\mathinner{|\psi_{ABC}\rangle}_{\lambda=\frac{1}{2}}\ =\ \frac{1}{2}\left(\mathinner{|000\rangle}+\mathinner{|011\rangle}-\right.$ $\sin{2\beta}\mathinner{|100\rangle}+$
@@
$H\ =\ \frac{1}{\sqrt{2}}\left[\begin{array}[]{cc}1&1\\ 1&-1\end{array}\right]\ .$

### v2

@@ Discussion
Finally, we should mention a proof that we skipped: The proof by Grassl, Beth, and Pellizzari shows that no ((3,2,3))(\!(3,2,3)\!) exists. As a consequence of, also no ((4,1,3))(\!(4,1,3)\!). This proof is rather intricate [GBP97, Theorem 5], relying on a sequence of arguments involving the Knill-Laflamme conditions, and the fact [GBP97, Lemma 3] that for every two–dimensional subspace of $\mathds{C}^{2}\otimes\mathds{C}^{2}$ there exists a basis that contains at least one product state. This leads to the conclusion that a ((3,2,3))(\!(3,2,3)\!) code can be factored into a tensor product, which contradicts that there does not exist a non-trivial code of length two.
@@
For future work, it will be interesting to see how much the semidefinite programming methods of Ref. [MNH24] will improve on the current bounds on quantum codes. We have used this method in shortened form in Proof 7: Lovász bound, but the full machinery has already proven the non-existence of $(\!(8,9,3)\!)_{2}$ and $(\!(10,5,4)\!)_{2}$ codes, both of which are not detected by the linear programming bounds [Rai98, HESG18]. A key aspect in this work is the state polynomial optimization framework, and we except that other non-linear problems can also be approached by this type of machinery [MH24, HMV24]. In these of applications, a key issue is typically the numerical precision of the infeasibility certificates involved. However, recent methods allow to extract rational infeasibility certificates, rendering these methods as good as any analytical proof [LdL24, PP08].
In the spirit of Refs. [CHLW14, CLW12] Another interesting research direction would be to develop an analytical or numerical machinery to systematically find new entropy and rank inequalities.
@@ Appendix A Three-qubit GHZ states
The proof is carried out by explicit construction of the local unitaries required to transform a given state to the standard GHZ form. Consider a three-qubit state $\mathinner{|\psi_{ABC}\rangle}$ with three maximally mixed one-body reduced density matrices. We start with a Schmidt decomposition with respect to the bipartite cut that contains qubit $A$,
@@
where $a$ and $b$ are complex numbers. This state can be substituted in $\mathinner{|\psi_{ABC}\rangle}$. If we now compute the reduced states $\rho_{B}$ and $\rho_{C}$ we find that the off-diagonal elements of $\rho_{B}$ - up to complex conjugation - are proportional to
(49) $\sqrt{\lambda}\ a^{*}\ +\ b\ \sqrt{1-\lambda}\ ,$
@@
To continue the proof, we must distinguish between two cases, (i) $\lambda\neq\frac{1}{2}$ and (ii) $\lambda=\frac{1}{2}$. (i) If $\lambda\neq\frac{1}{2}$ and $a,b\neq 0$ vanishing off-diagonal elements of the one-body reduced density matrices in Eq. (49) and (50) is not possible. Therefore it must hold that $a=b=0$. Hence for $\lambda\neq\frac{1}{2}$ we obtain for $\psi_{ABC}$,
@@
(55) $\left[\begin{array}[]{cc}\sqrt{\lambda}&-\sqrt{1-\lambda}\\ \sqrt{1-\lambda}&\sqrt{\lambda}\end{array}\right]_{A}\ \otimes\ \mathds{1}_{B}\ \otimes\ \mathds{1}_{C}$
@@
(ii) If instead $\lambda=\frac{1}{2}$ we have, according to Eq. (51), $b=a^{*}$ so that $\mathinner{|u_{1}\rangle}$ becomes
@@
[…]
@@
[…]